what will be the present value of my cost of 4yrs education when I turn 18yrs in 5yrs time if my current education cost is $12500 per year and is expected to remain constant for four years and I am 13yrs now.
Answer by Theo(13342)

you're 13 years old now.

you will be 18 years old in 5 years.

you will be paying 12,500 for 4 years starting 5 years from now.

your first payment will be 5 years from now.
your second payment will be 6 years from now.
your third payment will be 7 years from now.
your fourth payment will be 8 years from now.

the formula to use is f = p * (1 + r) ^ n

f is the future value
p is the present value
r is the interest rate per time period.
n is the number of time periods.

your future value is the same for all 4 years of college.

it is equal to 12,500

the formula for your first payment will be 12,500 = p * (1 + r) ^ 5
the formula for your second payment will be 12,500 = p * (1 + r) ^ 6
the formula for your third payment will be 12,500 = p * (1 + r) ^ 7
the formula for your fourth payment will be 12,500 = p * (1 + r) ^ 8

you would solve for p in each of these formulas to get:

p = 12,500 / (1 + r) ^ 5
p = 12,500 / (1 + r) ^ 6
p = 12,500 / (1 + r) ^ 7
p = 12,500 / (1 + r) ^ 8

r is whatever interest rate per year you were given.
for example, if the interest rate was given as 7%, then r would be equal to .07.

add up all the results and that's the present value of your 4 years of tuition costs starting 5 years from now.